5. Adult tickets for a play cost $5 and child tickets cost $1. If there were 20 people at a performance and the theater collected $92 from ticket sales, how many children attended the play?
A) 1 child
B) 2 children
C) 3 children
D) 18 children
no book Thanks so very much for your time!!!

Answer by ankor@dixie-net.com(22740)

Adult tickets for a play cost $5 and child tickets cost $1. If there were 20 people at a performance and the theater collected $92 from ticket sales, how many children attended the play?
:
Let A = number of adults, c = number of children
It says there were 20 people at the performance so the equation would be:
A + C = 20
:
It says adult tickets ae $5 and kids $1, the total was $92. The equation is:
5A + 1C = 92
:
If we take the 1st equation and rearrange it like this:
A + C = 20
C = 20 - A
:
Now substitute (20-A) for C in the 2nd equation and solve for A:
5A + 1C = 92
5A + 1(20 - A) = 92
5A - A + 20 = 92
4A = 92 - 20
4A = 72
A = 72/4
A = 18 adults
:
Find the number of kids
18 + C = 20
C = 20 - 18
C = 2 children
:
:
Check our solutions in the 2nd equation: 5(18) + 1(2) = 92